Cooling Tools, Inc. is currently producing 522 of small refrigerators per month but the company’s CEO plans to increase production at a rate of 8.82 percent per month until the firm is producing 7,471 of refrigerators per month. How many months will this take?

Round the answer to two decimal places.
Answer #1

FV = PV * (1 + r)n

7471 = 522 * (1 + 8.82%)n

14.3123 = 1.0882n

Taking log of both sides,

LN(14.3123) = n * LN(1.0882)

2.6611 = n * 0.0845

n = 31.48 months